OpenStudy (anonymous):

Can someone help me? c:

OpenStudy (anonymous):

Find k such that the line is tangent to the graph of the function. \[f(x)=k-x^{2}\]\[y=-4x+7\]

OpenStudy (anonymous):

First, find the derivative of f(x). You know how to do that?

OpenStudy (anonymous):

I have but the only problem i had was what would happen to k?

OpenStudy (♪chibiterasu):

They are two different equations. k never disappeared.

OpenStudy (anonymous):

If this is a Calculus class, I would find the derivative first.

OpenStudy (♪chibiterasu):

^^ I second this.

OpenStudy (anonymous):

At the point of tangency, both equations will have the same slope. That's why we want to find the derivative...so we can figure out at what point your 2 equations will have the same slope.

OpenStudy (anonymous):

f(x) = k - x^2 f'(x) = ______________

OpenStudy (anonymous):

k-2x

OpenStudy (anonymous):

That k is just a constant, so the derivative of that k is just 0...it goes away. Sok f ' (x) = -2x

OpenStudy (anonymous):

What is the slope of our line: y = -4x + 7 ?

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

So, set f'(x) equal to -4 and solve for x. What do you get?

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

Great, that's the x-coordinate of our point of tangency. To get the y coordinate, plug x=2 back into y = -4x + 7

OpenStudy (anonymous):

y=-1

OpenStudy (anonymous):

yes. So now take f(x) = k - x^2 and plucg in 2 for x and -1 for y (f(x) is the same as y) and solve for k and you're done!

OpenStudy (anonymous):

k=3

OpenStudy (anonymous):

That's right. Good job

OpenStudy (anonymous):

Oh man....Thank you for your help. :) You too Chibiterasu.

OpenStudy (anonymous):

A lot of people see that word "tangent" and don't think "oh, they have to have the same slope there" which is just the first derivative

OpenStudy (anonymous):

No problem.